# Scalar-induced gravitational waves from coherent initial states

### v1

@@ IV Spectral density of secondary GW
Notice that when $\alpha({\bm{k}})=\alpha(k)$, the condition to preserve reality of $\braket{{\mathcal{R}}({\bm{x}})}$, $\alpha(-{\bm{k}})=-\alpha^{\ast}({\bm{k}})$ simply becomes $\alpha(k)=-\alpha^{\ast}(k)$, enforcing $\alpha$ to be fully imaginary. Hence, preserving isotropy restores parity in $\braket{{\mathcal{R}}({\bm{x}})}$ and $h^{\lambda}_{\bm{k}}$. Further, the one-point function of $\braket{h^{\lambda}_{\bm{k}}(\eta)}$ simply vanishes in this case, due to azimuthal integration. Therefore, the effect of $\alpha(k)$ on secondary GW shall be solely through the partially reducible term $g^{\lambda_{1}\lambda_{2}}({\bm{k}}_{1},{\bm{k}}_{2})$ in the two-point function [cf. Eq. (34)].
@@
In order compute the inhomogeneous and homogeneous contributions in the two-point function, we adopt the technique used in [54]: we integrate the dimensionless combination of the two-point correlation $(k_{1}k_{2})^{3/2}\Braket{h^{\lambda_{1}}({\bm{k}}_{1})h^{\lambda_{2}}({\bm{k}}_{2})}/(2\pi^{2})$ over a small volume in Fourier space to arrive at an equivalent of the power spectrum, denoted as $\mathscr{P}_{{}_{\rm h}}^{\lambda_{1}\lambda_{2}}({\bm{k}}_{1},{\bm{k}}_{2},\eta)$. For an alternative yet equivalent derivation on spectral density, see Appendix A.
@@
To obtain $\mathscr{P}_{{}_{\rm h}}^{\lambda_{1}\lambda_{2}}({\bm{k}}_{1},{\bm{k}}_{2},\eta)$, with ${\bm{k}}_{2}=-{\bm{k}}_{1}$, we integrate over ${\bm{k}}_{2}$ over a volume of $k_{1}^{3}\Delta^{3}$ in the direction of $-\hat{\bm{k}}_{1}$, with $\Delta$ being a small dimensionless parameter. This gives us $\mathscr{P}_{{}_{\rm h}}^{\lambda_{1}\lambda_{2}}({\bm{k}}_{1},-{\bm{k}}_{1},\eta)$ to be
$\mathscr{P}_{{}_{\rm h}}^{\lambda_{1}\lambda_{2}}({\bm{k}},-{\bm{k}},\eta)$ $=\delta^{\lambda_{1}\lambda_{2}}\frac{k^{3}}{2\pi^{2}}\,\left[f^{\lambda_{1}}(k)+\right.\left.k^{3}\Delta^{3}\,g^{\lambda_{1}\lambda_{1}}({\bm{k}},-{\bm{k}})\right]\,,$ (38)
$=64\pi^{2}\,\delta^{\lambda_{1}\lambda_{2}}\,k^{3}\int\frac{{\rm d}^{3}\bm{q}}{(2\pi)^{3}}|Q^{\lambda}(\bm{k},\bm{q})|^{2}$
×$\times I^{2}(|\bm{k}-\bm{q}|,q,\eta)\frac{\mathcal{P}_{{}_{\mathrm{S}}}(q)}{q^{3}}\frac{\mathcal{P}_{{}_{\mathrm{S}}}(|{\bm{k}}-{\bm{q}}|)}{|\bm{k}-\bm{q}|^{3}}$
×$\times\left[1+2\Delta^{3}k^{3}\,|\alpha(|\bm{k}-\bm{q}|)|^{2}\right]\,.$ (39)
@@
$\mathscr{P}_{{}_{\rm h}}^{\lambda_{1}\lambda_{2}}({\bm{k}}_{1},{\bm{k}}_{2})$ $=$ $\left(\frac{k_{1}^{3}+k_{2}^{3}}{2}\right)\Delta^{3}\,\frac{(k_{1}k_{2})^{3/2}}{2\pi^{2}}\,g^{\lambda_{1}\lambda_{2}}({\bm{k}}_{1},{\bm{k}}_{2})\,,$ (41)
@@
×$\times\int\frac{{\rm d}^{3}\bm{q}}{(2\pi)^{3}}Q^{\lambda_{1}}(\bm{k}_{1},\bm{q})Q^{\lambda_{2}}(\bm{k}_{2},\bm{q})$
×$\times I(|\bm{k}_{1}-\bm{q}|,q,\tau)I(|\bm{k}_{2}+\bm{q}|,q,\tau)$
×$\times\frac{\mathcal{P}_{{}_{\mathrm{S}}}(q)}{q^{3}}\sqrt{\frac{\mathcal{P}_{{}_{\mathrm{S}}}(|{\bm{k}}_{1}-{\bm{q}}|)\mathcal{P}_{{}_{\mathrm{S}}}(|{\bm{k}}_{2}+{\bm{q}}|)}{|{\bm{k}}_{1}-{\bm{q}}|^{3}|{\bm{k}}_{2}+{\bm{q}}|^{3}}}$
×$\times\alpha(|{\bm{k}}_{1}-{\bm{q}}|)\alpha(|{\bm{k}}_{2}+{\bm{q}}|)\,.$
Note that in this term, there is no guarantee of $\delta^{\lambda_{1}\lambda_{2}}$. Hence, there can be cross-correlation between different polarizations in $\mathscr{P}_{{}_{\rm h}}^{\lambda_{1}\lambda_{2}}({\bm{k}}_{1},{\bm{k}}_{2})$ with ${\bm{k}}_{2}\neq-{\bm{k}}_{1}$. We shall examine the nature of correlations between different polarizations in different types of terms in a later section.
@@
$\Omega_{{}_{\rm GW}}(k)=\frac{k^{2}}{48a^{2}H^{2}}\sum_{\lambda}\mathcal{P}_{\mathrm{h}}^{\lambda}(k,\eta),$ (44)
@@
$\Omega_{{}_{\rm GW}}({\bm{k}}_{1},{\bm{k}}_{2})=\frac{k^{2}}{48a^{2}H^{2}}\sum_{\lambda}\mathscr{P}_{{}_{\rm h}}^{\lambda\lambda}(\bm{k}_{1},\bm{k}_{2},\eta).$ (45)
We may then evaluate the quantity today as [29, 30]
@@
where $h$ is defined in terms of Hubble parameter today $H_{0}$ as $H_{0}=100\,h\,{\rm km}\,{\rm s}^{-1}{\rm Mpc}^{-1}$.
IV.1 Parameterization of the primordial mean (live!)
@@
$\alpha(\bm{k})=-\frac{i}{k^{3/2}}\alpha_{0}F(k)\beta(\hat{\bm{k}}),$ (47)
where the dimensionless functions $F(k)$ and $\beta(\hat{\bm{k}})$ capture the magnitude and angular dependence of the primordial mean, respectively, while the dimensionless parameter $\alpha_{0}$ sets the overall amplitude. Since we preserve isotropy and focus just on inhomogeneity of $\alpha({\bm{k}})$ in this section, we set $\beta(\hat{\bm{k}})=1$.
We are specifically interested in generation of enhancement in $\Omega_{{}_{\rm GW}}$. Before proceeding with the numerical evaluation of $\Omega_{{}_{\rm GW}}$, we should note that the perturbativity of $\braket{{\mathcal{R}}}$ imposes a constraint on $\alpha_{0}$. Recall that we obtain $\mathcal{P}_{{}_{\mathrm{S}}}(k)\simeq H^{2}/(8\pi^{2}M_{{}_{\mathrm{Pl}}}^{2}\epsilon_{1})$ under slow-roll approximation. However, $\mathcal{P}_{{}_{\mathrm{S}}}(k)$ can deviate significantly from this estimate due to features in the potential and/or non-trivial background dynamics. If we assume $\mathcal{P}_{{}_{\mathrm{S}}}(k)$ to be scale-invariant $\mathcal{P}_{{}_{\mathrm{S}}}(k)=A_{\rm s}$ for simplicity and substitute Eq. (47) in Eq. (10), then $\braket{{\mathcal{R}}}<1\Rightarrow\alpha_{0}\sqrt{A_{\rm s}}<1$. Hence, for a given value of $A_{\rm s}$, we may not be able to set arbitrary values of $\alpha_{0}$, but ensure $\alpha_{0}<1/\sqrt{A_{\rm s}}$ While the amplitude $A_{\rm s}$ is well-constrained on large scales, it can be enhanced over small scales that remain largely unconstrained (see for instance [35, 68]). This shall tighten the bound on $\alpha_{0}$. On the other hand, if $A_{\rm s}$ remains small over small scales, then $\alpha_{0}$ can reach reasonably large values. We shall explore the interplay of these parameters in the amplitude of $\Omega_{{}_{\rm GW}}$.
@@
$\Omega_{{}_{\rm GW}}({\bm{k}},-{\bm{k}})=\frac{k^{2}}{48a^{2}H^{2}}\sum_{\lambda}\mathscr{P}_{{}_{\mathrm{h}}}^{\lambda\lambda}(\bm{k},-\bm{k},\eta).$ (49)
To evaluate $\mathscr{P}_{{}_{\mathrm{h}}}^{\lambda\lambda}(\bm{k},-\bm{k})$ numerically, it is convenient to align the wave vector $\bm{k}$ along the $\hat{\bm{q}}_{\bm{z}}$-direction. We also define the following dimensionless variables involved in the integration
@@
Recasting the integrals in terms of these variables and summing over the two polarization modes (+,×+$\times$+ , × or $R,L$), we get
$\sum_{\lambda}\mathscr{P}_{{}_{\mathrm{h}}}^{\lambda\lambda}(\bm{k},-\bm{k})$ $=8\int_{0}^{\infty}{\rm d}v\int_{-1}^{1}{\rm d}\mu\ \left(1-\mu^{2}\right)^{2}\frac{v^{3}}{u^{3}}$
×$\times\tilde{I}^{2}(u,v,x)\mathcal{P}_{{}_{\mathrm{S}}}(uk)\mathcal{P}_{{}_{\mathrm{S}}}(vk)$
×$\times\left[1+{\color[rgb]{1,0,0}}2\frac{\Delta^{3}\alpha_{0}^{2}}{2\pi\Delta_{f}^{2}u^{3}}\exp\left(-\frac{1}{\Delta_{f}^{2}}\ln^{2}\left(\frac{uk}{k_{0}}\right)\right)\right],$ (51)
@@
×$\times\left(I_{B}(u_{1},v_{1})I_{B}(u_{2},v_{2})+\pi^{2}I_{C}(u_{1},v_{1})I_{C}(u_{2},v_{2})\right),$ (52)
@@
$\tilde{I}_{B}(u,v)$ $\abs\displaystyle=-4uv+(u^{2}+v^{2}-3)\ln\abs$
$\tilde{I}_{C}(u,v)$ $=\left(u^{2}+v^{2}-3\right)\Theta(v+u-\sqrt{3}).$ (53b)
Owing to radiation domination, we set $aH=1/\eta$ in the above expression. Under these assumptions, we numerically evaluate $\sum_{\lambda}\mathscr{P}_{{}_{\mathrm{h}}}^{\lambda\lambda}({\bm{k}},-{\bm{k}})$ and hence $h^{2}\Omega_{{}_{\rm GW}}^{0}({\bm{k}},-{\bm{k}})$.
Figure 1 shows the enhancement in $\Omega_{{}_{\rm GW}}^{0}(\bm{k},-\bm{k})$ across a range of scales $k$ and $A_{\rm s}$. We set $\alpha_{0}=0.1/\sqrt{A_{\rm s}}$ to ensure perturbativity of $\braket{{\mathcal{R}}}$ discussed earlier. Figure 1(a) shows that the enhancement in $\Omega_{{}_{\rm GW}}^{0}$ due to the primordial mean is most pronounced for smaller values of $A_{\rm s}$. When $A_{\rm s}$ is comparable to its CMB-constrained value ($A_{\rm s}\sim 10^{-9}$), the spectral density peaks sharply as $k$ approaches $k_{0}$, with a maxima at $k\simeq 1.58\,k_{0}$. Post the peak, for $k>k_{0}$, the spectral density falls with a relatively flatter slope.
@@
$\Omega_{{}_{\rm GW}}({\bm{k}}_{1},{\bm{k}}_{2})=\frac{k^{2}}{48a^{2}H^{2}}\sum_{\lambda}\mathscr{P}_{{}_{\mathrm{h}}}^{\lambda\lambda}(\bm{k}_{1},\bm{k}_{2},\eta).$ (53bb)
@@
$\mu\equiv\hat{\bm{k}}_{1}\cdot\hat{\bm{q}}=-\hat{\bm{k}}_{2}\cdot\hat{\bm{q}}=\cos\theta_{,}\quad v\equiv\frac{q}{k},$ (53bc)
@@
×$\times\int_{0}^{\infty}{\rm d}v\int_{-1}^{1}{\rm d}\mu\ \frac{v^{3}}{u^{3}w^{3}}\left(1-\mu^{2}\right)^{2}$
×$\times\tilde{I}(u,v,x)\tilde{I}(w,v,x)$
×$\times\mathcal{P}_{{}_{\mathrm{S}}}(vk_{a})\sqrt{\mathcal{P}_{{}_{\mathrm{S}}}(uk_{a})\mathcal{P}_{{}_{\mathrm{S}}}(wk_{a})}$
×$\times\exp\left[-\frac{1}{2\Delta_{f}^{2}}\ln^{2}\left(\frac{uk_{a}}{k_{0}}\right)\right]$
×$\times\exp\left[-\frac{1}{2\Delta_{f}^{2}}\ln^{2}\left(\frac{wk_{a}}{k_{0}}\right)\right]\,,$ (53bd)
@@
$k_{a}(k_{1},k_{2})$ $\equiv\frac{k_{1}+k_{2}}{2},$ (53be)
$u(\mu,v)$ $\equiv\sqrt{v^{2}+\frac{k_{1}^{2}}{k_{a}^{2}}-2\frac{k_{1}}{k_{a}}v\mu},$ (53bf)
$w(\mu,v)$ $\equiv\sqrt{v^{2}+\frac{k_{2}^{2}}{k_{a}^{2}}-2\frac{k_{2}}{k_{a}}v\mu}.$ (53bg)
@@
Figure 2 shows the complete structure of $\Omega_{{}_{\rm GW}}({\bm{k}}_{1},{\bm{k}}_{2})$. As we can see, $\Omega_{{}_{\rm GW}}({\bm{k}}_{1},{\bm{k}}_{2})$ peaks on the diagonal $k_{1}=k_{2}$ and gradually vanishes at regions away from the diagonal line. The diagonal contribution arises from both the homogeneous and inhomogeneous components of $\mathscr{P}_{{}_{\rm h}}^{\lambda_{1}\lambda_{2}}({\bm{k}}_{1},{\bm{k}}_{2},\eta)$, whereas the off-diagonal contribution is solely from the inhomogeneous part arising from $\alpha(k)$. The spectral density peaks around the scale of inhomogeneity $k_{1}\approx k_{2}\approx k_{0}$, and decays away from $k_{0}$. It is interesting to note that the structure of $\Omega_{{}_{\rm GW}}({\bm{k}}_{1},{\bm{k}}_{2})$ is not symmetric about $k_{1}=-k_{2}$. This asymmetric structure is induced by the kernel of integration arising from factors of $\tilde{I}$ and polarization factors apart from $\alpha(k)$.

### v2

@@ IV Spectral density of secondary GW
Notice that when $\alpha({\bm{k}})=\alpha(k)$, the condition to preserve reality of $\braket{{\mathcal{R}}({\bm{x}})}$, $\alpha(-{\bm{k}})=-\alpha^{\ast}({\bm{k}})$ simply becomes $\alpha(k)=-\alpha^{\ast}(k)$, enforcing $\alpha$ to be fully imaginary. Hence, preserving isotropy restores parity in $\braket{{\mathcal{R}}({\bm{x}})}$ and $h^{\lambda}_{\bm{k}}$. Further, the one-point function of $\braket{h^{\lambda}_{\bm{k}}(\eta)}$ simply vanishes in this case, due to azimuthal integration. This can be understood by substituting the factor due to polarization tensor $Q^{\lambda}({\bm{k}},{\bm{q}})$ given in Eq. (22) or (23) in $\braket{h^{\lambda}_{\bm{k}}(\eta)}$ expressed in Eq. (29). With $\alpha(q)\alpha(|{\bm{k}}-{\bm{q}}|)$ being function of just $q$ and the polar angle $\theta$, the only dependence on the azimuthal angle $\phi$ arises in terms of $\exp(\pm 2i\phi)$ from $Q^{\lambda}({\bm{k}},{\bm{q}})$ in the integrand, which vanishes upon integration over $\phi$. Therefore, the effect of $\alpha(k)$ on secondary GW shall be solely through the partially reducible term $g^{\lambda_{1}\lambda_{2}}({\bm{k}}_{1},{\bm{k}}_{2}{,\eta})$ in the two-point function [cf. Eq. (34)].
@@
In order compute the inhomogeneous and homogeneous contributions in the two-point function, we adopt the technique used in [58]: we integrate the dimensionless combination of the two-point correlation $(k_{1}k_{2})^{3/2}\Braket{h^{\lambda_{1}}({\bm{k}}_{1})h^{\lambda_{2}}({\bm{k}}_{2})}/(2\pi^{2})$ over a small volume in Fourier space to arrive at an equivalent of the power spectrum, denoted as $\mathscr{P}_{{}_{\rm h}}^{\lambda_{1}\lambda_{2}}({\bm{k}}_{1},{\bm{k}}_{2},\eta)$.
@@
To obtain $\mathscr{P}_{{}_{\rm h}}^{\lambda_{1}\lambda_{2}}({\bm{k}}_{1},{\bm{k}}_{2},\eta)$, with ${\bm{k}}_{2}=-{\bm{k}}_{1}$, we integrate over ${\bm{k}}_{2}$ over a volume of $k_{1}^{3}\Delta^{3}$ in the direction of $\hat{\bm{k}}_{1}$, with $\Delta$ being a small dimensionless parameter. This gives us $\mathscr{P}_{{}_{\rm h}}^{\lambda_{1}\lambda_{2}}({\bm{k}}_{1},-{\bm{k}}_{1},\eta)$ to be
$\mathscr{P}_{{}_{\rm h}}^{\lambda_{1}\lambda_{2}}({\bm{k}},-{\bm{k}},\eta)$ $=\delta^{\lambda_{1}\lambda_{2}}\frac{k^{3}}{2\pi^{2}}\,\left[f^{\lambda_{1}}(k{,\eta})+\right$.
$\left.\qquad k^{3}\Delta^{3}\,g^{\lambda_{1}\lambda_{1}}({\bm{k}},-{\bm{k}}{,\eta})\right]\,$ (38)
The assumption in performing the integration is that $g({\bm{k}},-{\bm{k}})$ does not vary much in $k$-space over the small volume of $k^{3}\Delta^{3}$, so that it can be simply extracted out of the integration. Hence, the inhomogeneous part of the spectral density seemingly scales as $\Delta^{3}$. But we should note that this scaling is not valid for arbitrary values of $\Delta$. If $\Delta$ is larger than the characteristic scale of variation in $g({\bm{k}}_{1},{\bm{k}}_{2})$, the expression of Eq. (38) is not true anymore and one has to perform the integration over ${\bm{k}}_{2}$ accounting for the variation of $g({\bm{k}}_{1},{\bm{k}}_{2})$ within the integration volume. For our analysis, we shall work with value of $\Delta$ which satisfies this requirement. For an alternative yet equivalent derivation of spectral density, that does not rely on the smallness parameter $\Delta$, refer Appendix A.
Substituting the explicit forms of the functions in $\mathscr{P}_{{}_{\rm h}}^{\lambda_{1}\lambda_{2}}({\bm{k}},-{\bm{k}},\eta)$ given above, we get
$\mathscr{P}_{{}_{\rm h}}^{\lambda_{1}\lambda_{2}}({\bm{k}},-{\bm{k}},\eta)$ $=64\pi^{2}\,\delta^{\lambda_{1}\lambda_{2}}\,k^{3}\int\frac{{\rm d}^{3}\bm{q}}{(2\pi)^{3}}|Q^{\lambda}(\bm{k},\bm{q})|^{2}$
$\times I^{2}(|\bm{k}-\bm{q}|,q,\eta)\frac{\mathcal{P}_{{}_{\mathrm{S}}}(q)}{q^{3}}\frac{\mathcal{P}_{{}_{\mathrm{S}}}(|{\bm{k}}-{\bm{q}}|)}{|\bm{k}-\bm{q}|^{3}}$
$\times\left[1+2\Delta^{3}k^{3}\,|\alpha(|\bm{k}-\bm{q}|)|^{2}\right]\,.$ (39)
@@
$\mathscr{P}_{{}_{\rm h}}^{\lambda_{1}\lambda_{2}}({\bm{k}}_{1},{\bm{k}}_{2}{,\eta})$ $=$ $\left(\frac{k_{1}^{3}+k_{2}^{3}}{2}\right)\Delta^{3}\,\frac{(k_{1}k_{2})^{3/2}}{2\pi^{2}}$ (41)
$\times\,g^{\lambda_{1}\lambda_{2}}({\bm{k}}_{1},{\bm{k}}_{2}{,\eta})\,$
@@
$\times\int\frac{{\rm d}^{3}\bm{q}}{(2\pi)^{3}}Q^{\lambda_{1}}(\bm{k}_{1},\bm{q})Q^{\lambda_{2}}(\bm{k}_{2},\bm{q})$
$\times I(|\bm{k}_{1}-\bm{q}|,q,\eta)I(|\bm{k}_{2}+\bm{q}|,q,\eta)$
$\times\frac{\mathcal{P}_{{}_{\mathrm{S}}}(q)}{q^{3}}\sqrt{\frac{\mathcal{P}_{{}_{\mathrm{S}}}(|{\bm{k}}_{1}-{\bm{q}}|)\mathcal{P}_{{}_{\mathrm{S}}}(|{\bm{k}}_{2}+{\bm{q}}|)}{|{\bm{k}}_{1}-{\bm{q}}|^{3}|{\bm{k}}_{2}+{\bm{q}}|^{3}}}$
$\times\alpha(|{\bm{k}}_{1}-{\bm{q}}|)\alpha(|{\bm{k}}_{2}+{\bm{q}}|)\,.$
Note that in this term, there is no guarantee of $\delta^{\lambda_{1}\lambda_{2}}$. Hence, there can be cross-correlation between different polarizations in $\mathscr{P}_{{}_{\rm h}}^{\lambda_{1}\lambda_{2}}({\bm{k}}_{1},{\bm{k}}_{2}{,\eta})$ with ${\bm{k}}_{2}\neq-{\bm{k}}_{1}$. We shall examine the nature of correlations between different polarizations in different types of terms in a later section.
@@
$\Omega_{{}_{\rm GW}}(k{,\eta})=\frac{k^{2}}{48a^{2}H^{2}}\sum_{\lambda}\mathcal{P}_{\mathrm{h}}^{\lambda}(k,\eta),$ (44)
@@
$\Omega_{{}_{\rm GW}}({\bm{k}}_{1},{\bm{k}}_{2}{,\eta})=\frac{k^{2}}{48a^{2}H^{2}}\sum_{\lambda}\mathscr{P}_{{}_{\rm h}}^{\lambda\lambda}(\bm{k}_{1},\bm{k}_{2},\eta).$ (45)
Assuming that the secondary gravitational waves are induced in the radiation dominated era, we may then evaluate the quantity today as [31, 32]
@@
where $h$ is defined in terms of Hubble parameter today $H_{0}$ as $H_{0}=100\,h\,{\rm km}\,{\rm s}^{-1}{\rm Mpc}^{-1}$. Note that the spectral density becomes time-independent in this scenario, as we shall see later.
IV.1 Parameterization of the primordial mean
@@
[…]
@@
$\Omega_{{}_{\rm GW}}({\bm{k}},-{\bm{k}}{,\eta})=\frac{k^{2}}{48a^{2}H^{2}}\sum_{\lambda}\mathscr{P}_{{}_{\mathrm{h}}}^{\lambda\lambda}(\bm{k},-\bm{k},\eta).$ (49)
To evaluate $\mathscr{P}_{{}_{\mathrm{h}}}^{\lambda\lambda}(\bm{k},-\bm{k}{,\eta})$ numerically, it is convenient to align the wave vector $\bm{k}$ along the $\hat{\bm{q}}_{\bm{z}}$-direction. We also define the following dimensionless variables involved in the integration
@@
Recasting the integrals in terms of these variables and summing over the two polarization modes ($\times$ or $R,L$), we get
$\sum_{\lambda}\mathscr{P}_{{}_{\mathrm{h}}}^{\lambda\lambda}(\bm{k},-\bm{k}{,\eta})$ $=8\int_{0}^{\infty}{\rm d}v\int_{-1}^{1}{\rm d}\mu\ \left(1-\mu^{2}\right)^{2}\frac{v^{3}}{u^{3}}$
$\times\tilde{I}^{2}(u,v,x)\mathcal{P}_{{}_{\mathrm{S}}}(uk)\mathcal{P}_{{}_{\mathrm{S}}}(vk)$
$\times\Bigg[1+2\frac{\Delta^{3}\alpha_{0}^{2}}{2\pi\Delta_{f}^{2}u^{3}}\exp\left[-\frac{1}{\Delta_{f}^{2}}\ln^{2}\left(\frac{uk}{k_{0}}\right)\right]\Bigg]\,$$\qquad$ (51)
@@
$\times\left(I_{B}(u_{1},v_{1})I_{B}(u_{2},v_{2})+\pi^{2}I_{C}(u_{1},v_{1})I_{C}(u_{2},v_{2})\right),$ (52)
@@
$\tilde{I}_{B}(u,v)$ $=-4uv+(u^{2}+v^{2}-3)\ln\mathinner{\!\left\lvert\frac{3-(u+v)^{2}}{3-(u-v)^{2}}\right\rvert}$ (53b)
$\tilde{I}_{C}(u,v)$ $=\left(u^{2}+v^{2}-3\right)\Theta(v+u-\sqrt{3}).$ (53c)
Owing to radiation domination, we set $aH=1/\eta$ in the above expression. Consequently, the spectral density $\Omega_{{}_{\rm GW}}({\bm{k}},-{\bm{k}},\eta)$ becomes time-independent, and we drop the explicit $\eta$ dependence in what follows. Under these assumptions, we numerically evaluate $\sum_{\lambda}\mathscr{P}_{{}_{\mathrm{h}}}^{\lambda\lambda}({\bm{k}},-{\bm{k}},\eta)$ and hence $h^{2}\Omega_{{}_{\rm GW}}^{0}({\bm{k}},-{\bm{k}})$.
[…]
@@
$\Omega_{{}_{\rm GW}}({\bm{k}}_{1},{\bm{k}}_{2})=\frac{k^{2}}{48a^{2}H^{2}}\sum_{\lambda}\mathscr{P}_{{}_{\mathrm{h}}}^{\lambda\lambda}(\bm{k}_{1},\bm{k}_{2},\eta).$ (54)
@@
$\mu\equiv\hat{\bm{k}}_{1}\cdot\hat{\bm{q}}=-\hat{\bm{k}}_{2}\cdot\hat{\bm{q}}=\cos\theta_{,}\quad v\equiv\frac{q}{k},$ (55)
@@
$\times\int_{0}^{\infty}{\rm d}v\int_{-1}^{1}{\rm d}\mu\ \frac{v^{3}}{u^{3}w^{3}}\left(1-\mu^{2}\right)^{2}$
$\times\tilde{I}(u,v,x)\tilde{I}(w,v,x)$
$\times\mathcal{P}_{{}_{\mathrm{S}}}(vk_{a})\sqrt{\mathcal{P}_{{}_{\mathrm{S}}}(uk_{a})\mathcal{P}_{{}_{\mathrm{S}}}(wk_{a})}$
$\times\exp\left[-\frac{1}{2\Delta_{f}^{2}}\ln^{2}\left(\frac{uk_{a}}{k_{0}}\right)\right]$
$\times\exp\left[-\frac{1}{2\Delta_{f}^{2}}\ln^{2}\left(\frac{wk_{a}}{k_{0}}\right)\right]\,$, (56)
@@
$k_{a}(k_{1},k_{2})$ $\equiv\frac{k_{1}+k_{2}}{2},$ (57)
$u(\mu,v)$ $\equiv\sqrt{v^{2}+\frac{k_{1}^{2}}{k_{a}^{2}}-2\frac{k_{1}}{k_{a}}v\mu},$ (58)
$w(\mu,v)$ $\equiv\sqrt{v^{2}+\frac{k_{2}^{2}}{k_{a}^{2}}-2\frac{k_{2}}{k_{a}}v\mu}.$ (59)
@@
Figure 2 shows the complete structure of $\Omega_{{}_{\rm GW}}({\bm{k}}_{1},{\bm{k}}_{2})$. As we can see, $\Omega_{{}_{\rm GW}}({\bm{k}}_{1},{\bm{k}}_{2})$ peaks on the diagonal $k_{1}=k_{2}$ and gradually vanishes at regions away from the diagonal line. The diagonal contribution arises from both the homogeneous and inhomogeneous components of $\mathscr{P}_{{}_{\rm h}}^{\lambda_{1}\lambda_{2}}({\bm{k}}_{1},{\bm{k}}_{2},\eta)$, whereas the off-diagonal contribution is solely from the inhomogeneous part arising from $\alpha(k)$. The spectral density peaks around the scale of inhomogeneity $k_{1}\approx k_{2}\approx k_{0}$, and decays away from $k_{0}$. It is interesting to note that the structure of $\Omega_{{}_{\rm GW}}({\bm{k}}_{1},{\bm{k}}_{2})$ is not symmetric about the line $\ln(k_{2}/(1.58\,k_{0}))=-\ln(k_{1}/(1.58\,k_{0}))$ in Fig. 2. This asymmetric structure is induced by the kernel of integration arising from factors of $\tilde{I}$ and polarization factors apart from $\alpha(k)$.